The sum of 6 times petras age and 8 times kathys age is 162. Kathy is 1 year more that twice as old at petra is. Find each of there ages
Answer by addingup(3677)

So the problem gives us two equalities, so we'll have two equations:
6p+8k= 162 (first equation)
------------------------------------
The second:
k= 2p+1 Let's substitute for k in the first equation with this value for k
-------------
6p+8(2p+1)= 162 We are done, all we have left to do is use our calculators:
6p+16p+8= 162 Add p on left and subtract 8 from both sides:
22p= 156
p= 7 This is how old Petra is. And the problem says that:
k= 2p+1
k=2(7)+1
k= 15 Kathy is 15
Proof - The problem says that:
6p+8k= 162
6(7)+8(15)= 162
42+120= 162
162= 162 We have the correct answer